OpenStudy (anonymous):

find four consecutive integers such that the product of the two largest is 46 more than the product of the two smallest ?

OpenStudy (anonymous):

let the 4 integers be x, x+1, x+2,x+3 then (x+2)(x+3) = y + 46 x(x+1) = y now its a matter of solving these two equations

OpenStudy (anonymous):

next step would be (x+2)(x+3) = x(x+1) + 46

OpenStudy (anonymous):

or the next step is enter those 2 things into a graphing calculator and find the intersection, which is x = 10, y = 110

OpenStudy (anonymous):

expanding x^2 + 5x + 6 = x^2 + x + 46 5x - x = 46 - 6 4x = 40 x = 10 - the rest is easy

OpenStudy (anonymous):

do u follow that ok?

OpenStudy (anonymous):

im doing it

OpenStudy (anonymous):

if you're feeling extra awesome and redundant, you could even put it into an augmented matrix and solve it that way, lol

OpenStudy (anonymous):

yep